"in a chemistry class, 7 liters of a 4% silver iodide solution must be mixed with a 10% solution to get a 6% solution. How many liters of the 10% solution are needed?"
Answer by stanbon(48558)

in a chemistry class, 7 liters of a 4% silver iodide solution must be mixed with a 10% solution to get a 6% solution. How many liters of the 10% solution are needed
--------------------
Equation:
iodide + iodide = iodide
0.04*7 + 0.10x = 0.06(7+x)
---
Multiply thru by 100 to get:
4*7 + 10x = 6(7+x)
4*7 + 10x = 6*7 + 6x
4x = 2*7
x = 3.5 liters (amount of 10% solution needed for the mixture)
=================================
Cheers,
Stan H.